Given
h(x) = 2x + 4
and
f(x) =
1/2
x + 9,
evaluate the composite function.
f[h(x)]

Answer by fractalier(6550)

To find f(h(x)) we plug h(x) in for x in the f(x) function and get
f(h(x)) = f(2x + 4) = (1/2)(2x+4) + 9 = x + 2 + 9 = x + 11